A 20 A Given the following, find PS and CS. P 500-50 5. FC 0 MC 300 D A Des M
Answer #1

In long run MC is the supply curve as the expansion in production cause no change in total cost of producing good.

P = 500 - 5Q ......(1)

Thus at equilibrium, P = MC

500 - 5Q = 300

200 = 5Q

Q = 40

Thus the consumer surplus is the shaded area whose area is (1/2) * 40 * (500 - 300) = 4000 while the producer surplus is zero as the supply curve is horizontal. Price 300 AS - - - - no 100 Quentity wy